OpenStudy (anonymous):

i am so stumped on this one Simplify 10(2x + 3) - 20. 20x + 10 20x - 17 20x - 10

OpenStudy (uri):

Multiply the bracket with 10 first What do you get?

OpenStudy (uri):

10(2x + 3) ?

OpenStudy (anonymous):

i dont get it

OpenStudy (phi):

you should learn the "distributive law" if you have 10 times ( 2x +3) that means you have 10 times each thing in side the parens think of the parens as a package, and you have 10 identical packages

OpenStudy (phi):

so how many 2 x's do you have and how many 3 's do you have

OpenStudy (anonymous):

ten

OpenStudy (anonymous):

each

OpenStudy (phi):

yes, and you write it 10(2x+3) = 10* 2x + 10*3 10*3 is easy, right ? 10 * 2x means 10*2*x you can simplify the 10*2 part

OpenStudy (uri):

10(2x+3) = 10* 2x + 10*3 =20x+30

OpenStudy (uri):

Now,20x+30-20 What is 30-20? @dirtydan667

OpenStudy (anonymous):

10

OpenStudy (uri):

So 20x+10 ans